### v1

# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

Abstract.
Shuffling has been shown to amplify differential privacy guarantees, offering a stronger privacy-utility trade-off. To characterize and compute this amplification, two fundamental analytical frameworks have been proposed: the privacy blanket by Balle et al. (CRYPTO 2019) and the clone paradigm (including both the standard clone and stronger clone) by Feldman et al. (FOCS 2021, SODA 2023). All these methods rely on decomposing local randomizers.
In this work, we introduce a unified analysis framework—the general clone paradigm—which encompasses all possible decompositions. We identify the optimal decomposition within the general clone paradigm. Moreover, we develop a simple and efficient algorithm to compute the exact value of the optimal privacy amplification bounds via Fast Fourier Transform. Experimental results demonstrate that the computed upper bounds for privacy amplification closely approximate the lower bounds, highlighting the tightness of our approach. Finally, using our algorithm, we conduct the first systematic analysis of the joint composition of LDP protocols in the shuffle model.
Differential privacy, shuffle model, general clone, fast fourier transform, joint composition
copyright: noneccs: Security and privacy Privacy-preserving protocols
1. Introduction
Differential Privacy (DP) has become a foundational framework for safeguarding individual privacy while enabling meaningful data analysis (Dwork, 2006). In real-world applications, Local Differential Privacy (LDP) is widely adopted as it eliminates the need for a trusted curator by applying noise to each user’s data before aggregation (Cormode et al., 2018; Ding et al., 2017; Yang et al., 2024a; Team, 2017). However, this decentralized approach often results in significant utility loss due to excessive noise.
To address this trade-off, the shuffle model introduces a trusted shuffler between users and the aggregator (Cheu et al., 2019; Erlingsson et al., 2019; Bittau et al., 2017). The shuffler permutes the locally perturbed data, breaking the link between individual users and their submitted values. Shuffle DP retains the trust-free nature of LDP while significantly improving the privacy-utility trade-off, making it a promising model for real-world deployment (Wang et al., 2020; Imola et al., 2022; Cheu and Zhilyaev, 2022). For instance, privacy amplification by shuffling was used in Apple and Google’s Exposure Notification Privacy-preserving Analytics (Apple and Google, 2021).
The amplification effect in the shuffle model means that when each of $n$ clients randomizes their data using an $\epsilon_{0}$-LDP mechanism, the shuffled reports satisfy $(\epsilon(\epsilon_{0},\delta,n),\delta)$-DP, where $\epsilon(\epsilon_{0},\delta,n)\ll\epsilon_{0}$ for sufficiently large $n$ and $\frac{1}{\delta}$ (Feldman et al., 2023a). A central theoretical challenge is to bound and compute this privacy amplification effect. A tighter bound enables the use of a larger $\epsilon_{0}$, achieving $(\epsilon,\delta)$-DP after shuffling and thereby improving utility.
Among the various strategies proposed to analyze this effect (Cheu et al., 2019; Erlingsson et al., 2019; Balle et al., 2020), decomposition-based approaches have demonstrated the strongest performance (Balle et al., 2019; Feldman et al., 2022, 2023a). Two prominent frameworks are the privacy blanket by Balle et al. (Balle et al., 2019) and the clone paradigm by Feldman et al. (Feldman et al., 2022, 2023a), which includes the standard clone and the stronger clone. These approaches rely on decomposing the probability distributions induced by a local randomizer under different inputs. A decomposition naturally leads to a reduction, which yields an upper bound on the privacy amplification achieved by shuffling.
The privacy blanket provides a tailored bound for each specific local randomizer through a decomposition, and further scaling yields a looser but computable bound. It also derives a computable bound for generic randomizers in a similar way. The standard clone adopts a simple, unified decomposition, yielding a generic bound for all $\epsilon_{0}$-DP local randomizers (Feldman et al., 2022). The generic bound can be computed precisely using a numerical algorithm, and is better than the privacy blanket’s computable bound for generic randomizers.
The stronger clone was proposed with a more refined decomposition, which was expected to provide a new bound for generic local randomizers and bounds for specific local randomizers (Feldman et al., 2023a). Unfortunately, a critical flaw was later found in the proof’s core lemma. A corrected version was published on arXiv (Feldman et al., 2023b), which showed that the original general bounds only hold for a restricted class of local randomizers. The original specific bounds were replaced with a version that is much weaker and cannot be computed efficiently. Although the authors conjecture that the original results may still hold in general, this remains unproven. Moreover, the flawed result has been propagated in follow-up works (Wang et al., 2024; Chen et al., 2024; Yang et al., 2024b; Wang et al., 2025).
In this paper, we provide the optimal bound for each specific randomizer among all possible decompositions, and meanwhile provide a numerical algorithm to efficiently and precisely calculate the bound.
To identify the optimal bound, we propose a unified analysis framework—the general clone paradigm-which encompasses all possible decompositions, and further show the best decomposition is the one used by the privacy blanket. However, the decomposition bound of the privacy blanket in its original form cannot be directly computed, which is why it was further scaled with Hoeffding’s and Bennett’s inequalities. Fortunately, we represent the decomposition bound in a simple form and propose a numerical algorithm using Fast Fourier Transform (FFT) to efficiently compute the bound.
With these results, we achieve the best-known bounds obtainable from decomposition-based methods. Experiments show that our computed upper bounds closely match empirical lower bounds, demonstrating the tightness of our analysis. Additionally, we discuss a potential direction to move beyond decomposition methods: identifying the exact amplification without bounding. While this approach is intuitively appealing, it currently lacks the necessary theoretical tools and remains an open problem.
Finally, we conduct the first systematic analysis of joint composition in the shuffle model using our algorithm. In classical DP, $k$-fold composition refers to applying $k$ independent mechanisms to the same dataset:
$\mathcal{M}_{kFold}(D)=(\mathcal{M}_{1}(D),\mathcal{M}_{2}(D),\dots,\mathcal{M}_{k}(D)).$
In contrast, our notion of joint composition applies independent mechanisms to different datasets:
$\mathcal{M}_{joint}(D_{1},D_{2},\dots,D_{k})=(\mathcal{M}_{1}(D_{1}),\mathcal{M}_{2}(D_{2}),\dots,\mathcal{M}_{k}(D_{k})).$
Joint composition is widely used in LDP applications such as joint distribution estimation and heavy hitter detection (Domingo-Ferrer and Soria-Comas, 2022; Ren et al., 2018; Bassily and Smith, 2015; Kikuchi, 2022). For example, when each user’s data contains $d$ attributes, applying an $\frac{\epsilon}{d}$-LDP mechanism to each attribute ensures overall $\epsilon$-LDP while preserving inter-attribute correlations.
While existing studies have analyzed $k$-fold composition in the shuffle model (Koskela et al., 2021, 2020), we focus on the case where the local mechanism itself is a joint composition. Our experimental results show that existing methods yield relatively loose bounds in this setting, whereas our algorithm computes significantly tighter results by leveraging the optimal bounds.
Our contributions can be summarized as follows:
  • We propose the general clone paradigm, which subsumes all decomposition-based methods, and identify the optimal bound it can provide for a specific local randomizer.
  • We provide an efficient numerical algorithm for computing the optimal bounds via FFT.
  • We perform the first systematic analysis of joint composition in the shuffle model, achieving significantly tighter amplification bounds.
2. Preliminaries
Differential privacy is a privacy-preserving framework for randomized algorithms. Intuitively, an algorithm is differentially private if the output distribution does not change significantly when a single individual’s data is modified. This ensures that the output does not reveal substantial information about any individual in the dataset. The hockey-stick divergence is commonly used to define $(\epsilon,\delta)$-DP.
Definition 0 (Hockey-Stick Divergence).
The hockey-stick divergence between two random variables $P$ and $Q$ is defined as:
$D_{\alpha}(P\parallel Q)=\int\max\{0,P(x)-\alpha Q(x)\}\,\mathrm{d}x,$
where we use the notation $P$ and $Q$ to refer to both the random variables and their probability density functions.
We say that $P$ and $Q$ are $(\epsilon,\delta)$-indistinguishable if:
$\max\{D_{e^{\epsilon}}(P\parallel Q),D_{e^{\epsilon}}(Q\parallel P)\}\leq\delta.$
If two datasets $X^{0}$ and $X^{1}$ have the same size and differ only by the data of a single individual, they are referred to as neighboring datasets (denoted by $X^{0}\simeq X^{1}$).
Definition 0 (Differential Privacy).
An algorithm $\mathcal{R}:\mathbb{X}^{n}\to\mathbb{Z}$ satisfies $(\epsilon,\delta)$-differential privacy if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, $\mathcal{R}(X)$ and $\mathcal{R}(X^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
Definition 0 (Local Differential Privacy).
An algorithm $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ satisfies local $(\epsilon,\delta)$-differential privacy if for all $x,x^{\prime}\in\mathbb{X}$, $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
Here, $\epsilon$ is referred to as the privacy budget, which controls the privacy loss, while $\delta$ allows for a small probability of failure. When $\delta=0$, the mechanism is also called $\epsilon$-DP.
Following conventions in the shuffle model based on randomize-then-shuffle (Balle et al., 2019; Cheu et al., 2019), we define a single-message protocol $\mathcal{P}$ in the shuffle model as a pair of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{O}$. We call $\mathcal{R}$ the local randomizer, $\mathbb{Y}$ the message space of the protocol, $\mathcal{A}$ the analyzer, and $\mathbb{O}$ the output space.
The overall protocol implements a mechanism $\mathcal{P}:\mathbb{X}^{n}\to\mathbb{O}$ as follows: Each user $i$ holds a data record $x_{i}$, to which they apply the local randomizer to obtain a message $y_{i}=\mathcal{R}(x_{i})$. The messages $y_{i}$ are then shuffled and submitted to the analyzer. Let $\mathcal{S}(y_{1},\dots,y_{n})$ denote the random shuffling step, where $\mathcal{S}:\mathbb{Y}^{n}\to\mathbb{Y}^{n}$ is a shuffler that applies a random permutation to its inputs.
In summary, the output of $\mathcal{P}(x_{1},\dots,x_{n})$ is given by
$\mathcal{A}\circ\mathcal{S}\circ\mathcal{R}^{n}(\boldsymbol{x})=\mathcal{A}(\mathcal{S}(\mathcal{R}(x_{1}),\dots,\mathcal{R}(x_{n}))).$
Definition 0 (Differential Privacy in the Shuffle Model).
A protocol $\mathcal{P}=(\mathcal{R},\mathcal{A})$ satisfies $(\epsilon,\delta)$-differential privacy in the shuffle model if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, the distributions $\mathcal{S}\circ\mathcal{R}^{n}(X)$ and $\mathcal{S}\circ\mathcal{R}^{n}(X^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
3. Review of Existing Analysis Techniques
In this section, we review existing analysis techniques for studying privacy amplification in the shuffle model. We begin by introducing the standard clone paradigm due to its simplicity (Feldman et al., 2022). Next, we restate the privacy blanket framework, using consistent terminology with the former (Balle et al., 2019). On one hand, this clarifies misconceptions about the privacy blanket framework found in subsequent literature (Koskela et al., 2021). On the other hand, it aids in identifying the intrinsic connection between the two approaches, which will be explored in Section 4.
We then discuss the subsequent attempts to extend the clone paradigm, specifically the stronger clone. We show the vision and failure of both the original and corrected versions of the stronger clone.
3.1. Standard clone
The intuition behind the standard clone paradigm is as follows (Feldman et al., 2022): Suppose that $X^{0}$ and $X^{1}$ are neighbouring databases that differ on the first datapoint, $x_{1}^{0}\neq x_{1}^{1}$. A key observation is that for any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}$ and data point $x$, $\mathcal{R}(x)$ can be seen as sampling from the same distribution as $\mathcal{R}(x_{1}^{0})$ with probability at least $e^{-\varepsilon_{0}}/2$ and sampling from the same distribution as $\mathcal{R}(x_{1}^{1})$ with probability at least $e^{-\varepsilon_{0}}/2$. That is, with probability $e^{-\varepsilon_{0}}$ each data point can create a clone of the output of $\mathcal{R}(x_{1}^{0})$ or a clone of $\mathcal{R}(x_{1}^{1})$ with equal probability. Thus $n-1$ data elements effectively produce a random number of clones of both $x_{1}^{0}$ and $x_{1}^{1}$, making it more challenging to distinguish whether the original dataset contains $x_{1}^{0}$ or $x_{1}^{1}$ as its first element.
Due to the $\epsilon_{0}$-DP property of the local randomizer $\mathcal{R}$, we have the following inequality:
$\forall x_{i}\in\mathbb{X},\forall y\in\mathbb{Y}:\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\epsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{0})=y]$
$\wedge\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\epsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{1})=y].$
Therefore, the local randomizer $\mathcal{R}$ on any input $x_{i}$ can be decomposed into a mixture of $\mathcal{R}(x_{1}^{0}),\mathcal{R}(x_{1}^{1})$ and some “left-over” distribution $\text{LO}(x_{i})$ such that
$\mathcal{R}(x_{i})=\frac{1}{2e^{\epsilon_{0}}}\mathcal{R}(x_{1}^{0})+\frac{1}{2e^{\epsilon_{0}}}\mathcal{R}(x_{1}^{1})+(1-\frac{1}{e^{\epsilon_{0}}})\text{LO}(x_{i}).$
Let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ denote the shuffling of $\mathcal{R}$. To compute the privacy amplification provided by the shuffle model, we need to compute $D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1}))$ for a given $\epsilon$. The exact computation is computationally complex, so the researchers seek an upper bound for it. A key property is that hockey-stick divergence satisfies the data processing inequality.
Property 1 (Data Processing Inequality).
For all distributions $P$ and $Q$ defined on a set $S$ and (possibly randomized) functions $f:S\to S^{\prime}$,
$D_{\alpha}\big{(}f(P)||f(Q)\big{)}\leq D_{\alpha}(P||Q).$
If we can find two probability distributions $P_{0}$ and $P_{1}$ along with a post-processing function $f$ such that $f(P_{0})=\mathcal{M}_{S}(X^{0})$ and $f(P_{1})=\mathcal{M}_{S}(X^{1})$, then it follows that $D_{e^{\epsilon}}(P_{0},P_{1})$ is an upper bound for $D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1}))$. We refer to $(P_{0},P_{1})$ as a reduction pair. Different analysis techniques construct different reduction pairs. We first present an intuitive construction of the reduction pair within the standard clone framework, followed by the formal construction.
Definition 0 (Standard Clone Reduction Pair (Intuitive) (Feldman et al., 2022)).
Define random variables $A_{0},A_{1}$ and $A_{2}$ as follows:
$A_{0}=\begin{cases}0&\text{w.p. }1\\ 1&\text{w.p. }0\\ 2&\text{w.p. }0\end{cases},\ A_{1}=\begin{cases}0&\text{w.p. }0\\ 1&\text{w.p. }1\\ 2&\text{w.p. }0\end{cases},\text{ and }A_{2}=\begin{cases}0&\text{w.p. }\frac{1}{2e^{\epsilon_{0}}}\\ 1&\text{w.p. }\frac{1}{2e^{\epsilon_{0}}}\\ 2&\text{w.p. }1-\frac{1}{e^{\epsilon_{0}}}\end{cases}$
To obtain a sample from $P_{0}$ (or $P_{1}$), sample one copy from $A_{0}$ (or $A_{1}$) and $n-1$ copies of $A_{2}$, the output $(n_{0},n_{1})$ where $n_{0}$ is the total number of 0s and $n_{1}$ is the total number of 1s. Equivalently,
$C\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon_{0}}}),\ A\sim\text{Bin}(C,\frac{1}{2}).$
$P_{0}=(A+1,C-A),\ P_{1}=(A,C-A+1).$
The corresponding post-processing function $f_{1}$ is shown in the Algorithm 1.
Algorithm 1 Post-processing function of standard clone (Feldman et al., 2022), $f_{1}$
$x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}$; $y\in\{0,1,2\}^{n}$
$J\leftarrow\emptyset$
for $i=1,\dots,n$ do
     if $y_{i}=2$ then
         Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
         $J\leftarrow J\cup\{j\}$
     end if
     Sample $z_{i}$ from
      $\begin{cases}\mathcal{R}(x_{1}^{0})&\text{if }y_{i}=0;\\ \mathcal{R}(x_{1}^{1})&\text{if }y_{i}=1;\\ \text{LO}(x_{j})&\text{if }y_{i}=2.\end{cases}$
end for
return $z_{1},\dots,z_{n}$
An additional observation is that if $\mathcal{R}$ is $\epsilon_{0}$-DP, then $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are similar, hence privacy is further amplified (Feldman et al., 2022). The similarity is characterized by the following lemma:
Lemma 2 ((Kairouz et al., 2017)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be an $\epsilon_{0}$-DP local randomizer and $x_{0},x_{1}\in\mathbb{X}$. Then there exists two probability distributions $\mathcal{Q}_{0},\mathcal{Q}_{1}$ such that
$\mathcal{R}(x_{0})=\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{0}+\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}$
and
$\mathcal{R}(x_{1})=\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{0}+\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}.$
With the help of Lemma 2, (Feldman et al., 2022) gives the following decomposition for generic local randomizers:
$\mathcal{R}(x_{1}^{0})$ $=\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\mathcal{R}(x_{1}^{1})$ $=\frac{1}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{0}+\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}\mathcal{Q}_{1}^{1},$
$\forall i\in[2,n]:\mathcal{R}(x_{i})$ $=\frac{1}{2e^{\epsilon_{0}}}\mathcal{Q}_{1}^{0}+\frac{1}{2e^{\epsilon_{0}}}\mathcal{Q}_{1}^{1}+(1-\frac{1}{e^{\epsilon_{0}}})\text{LO}(x_{i}).$
This decomposition leads to the formal reduction of the standard clone:
Theorem 3 (Standard Clone Reduction (Feldman et al., 2022)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\epsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}^{C},P_{1}^{C}\big{)}$
where $P_{0}^{C},\ P_{1}^{C}$ are defined as below (with “C” denoting “standard Clone”) :
$C\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon_{0}}}),\quad A\sim\text{Bin}(C,\frac{1}{2}),\quad\text{and }\quad\Delta\sim\text{Bern}(\frac{e^{\epsilon_{0}}}{e^{\epsilon_{0}}+1}).$
$P_{0}^{C}=(A+\Delta,C-A+1-\Delta),\quad P_{1}^{C}=(A+1-\Delta,C-A+\Delta).$
Bern($p$) represents a Bernoulli random variable with bias $p$.
Proof.
We can construct a post-processing function from $(P_{0}^{C},P_{1}^{C})$ to $\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$, which is similar to Algorithm 1. The only difference is that $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ are replaced by $\mathcal{Q}_{1}^{0}$ and $\mathcal{Q}_{1}^{1}$, respectively. ∎
3.2. Privacy Blanket Framework
The decomposition of the standard clone paradigm is based on the projections of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$, using them as reference points and projecting $\mathcal{R}(x_{i})$ onto these bases (Feldman et al., 2022). In contrast, the decomposition provided by the privacy blanket framework first computes the “common part” of all $\mathcal{R}(x)$ (Balle et al., 2019):
$\omega(y)=\inf_{x\in\mathbb{X}}\mathcal{R}(x)(y)/\gamma,$
where $\mathcal{R}(x)(y)$ is the probability density of $\mathcal{R}(x)$ at point $y$, and $\gamma$ is a normalization factor:
$\gamma=\int\inf_{x}\mathcal{R}(x)(y)\,\mathrm{d}y.$
Here, $\omega$ and $\gamma$ are referred to as the privacy blanket distribution and the total variation similarity of the local randomizer $\mathcal{R}$ (Balle et al., 2019). Each $\mathcal{R}(x)$ can then be decomposed as:
$\mathcal{R}(x)=(1-\gamma)\text{LO}(x)+\gamma\omega,$
where $\text{LO}(x)$ represents the ”left-over” distribution.
In other words, the execution of each $\mathcal{R}(x_{i})$ can be viewed as first sampling a random variable $b_{i}\sim\text{Bern}(\gamma)$. If $b_{i}=1$, a sample is drawn from $\omega$ and returned; otherwise, a sample is drawn from $\text{LO}(x_{i})$.
The original proof is formulated using the terminology of the “View” of the server. However, we observe that some subsequent works have misinterpreted its meaning (Koskela et al., 2021). To clarify, we restate the privacy blanket technique using the following notation: probability distributions $P_{0}^{B}$ and $P_{1}^{B}$ (with “B” denoting “Blanket”), along with a post-processing function $f^{B}$.
Definition 0 (Privacy Blanket Reduction Pair (Balle et al., 2019) (Restated)).
Let $\boldsymbol{x}_{-1}=(x_{2},x_{3},\dots,x_{n})$ with the inputs from the last $n-1$ users, $\boldsymbol{y}^{a}=(y_{1}^{a},y_{2},\dots,y_{n})$ where $y_{i}\sim R(x_{i})$ is the output of the i-th user, $a$ indicates that the input of the first user is $x_{1}^{a}$, $a\in\{0,1\}$. Let $\boldsymbol{b}=(b_{2},b_{3},\dots,b_{n})$ be binary values indicating which users sample from the privacy blanket distribution. A multiset $Y_{\boldsymbol{b}}^{a}=\mathcal{S}(y_{1}^{a}\cup\{y_{i}|b_{i}=1\})$.
Observe that the distribution of $Y_{\boldsymbol{b}}^{a}$ depends only on $|\boldsymbol{b}|$ rather than $\boldsymbol{b}$, where $|\boldsymbol{b}|$ represents the number of 1 in $\boldsymbol{b}$. We can rewrite it as $Y_{|\boldsymbol{b}|}^{a}=\mathcal{S}({y_{1}}\cup\{y_{i}|y_{i}\sim\omega,i=1,2,\dots,|\boldsymbol{b}|\})$. Then $P_{0}^{B}$ and $P_{1}^{B}$ are defined below:
$P_{0}^{B}=(|\boldsymbol{b}|,Y^{0}_{|\boldsymbol{b}|}),$
$P_{1}^{B}=(|\boldsymbol{b}|,Y^{1}_{|\boldsymbol{b}|}),$
where $|\boldsymbol{b}|\sim\text{Bin}(n-1,\gamma)$.
Theorem 5 (Privacy Blanket Reduction (Balle et al., 2019) (Restated)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a $\epsilon_{0}$-DP local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. For $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$, we have
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}^{B},P_{1}^{B}\big{)}.$
Proof.
The corresponding post-processing function $f^{B}$ is shown in Algorithm 2. The core idea of the post-processing function $f^{B}$ is that, given $Y_{|\boldsymbol{b}|}^{a}$, it suffices to sample from the left-over distributions of a randomly selected subset of $n-1-|\boldsymbol{b}|$ users and mix the results accordingly. ∎
Algorithm 2 Post-processing function of privacy blanket, $f^{B}$
$x_{2},\dots,x_{n}$; $|\boldsymbol{b}|\in\{0,1,\dots,n-1\},Y^{a}_{|\boldsymbol{b}|}$
$J\leftarrow\emptyset$
$S\leftarrow\emptyset$
for $loop=1,\dots,n-1-|\boldsymbol{b}|$ do
     Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
     $s\leftarrow_{\text{LO}(x_{j})}\mathbb{Y}$triangleright Sample from $\text{LO}(x_{j})$
     $J\leftarrow J\cup\{j\}$
     $S\leftarrow S\cup\{s\}$
end for
return $Y^{a}_{|\boldsymbol{b}|}\cup S$
In Appendix A, we point out and correct the misunderstanding of the privacy blanket framework in (Koskela et al., 2021).
3.3. Vision and failure of stronger clone
The stronger clone is expected to improve the probability of producing a “clone” from $\frac{1}{2e^{\epsilon_{0}}}$ to $\frac{1}{e^{\epsilon_{0}}+1}$. For $\epsilon_{0}>1$, this results in approximately a factor of 2 improvement in the expected number of “clones” (Feldman et al., 2023a). This improvement is anticipated to be achieved through a more refined analysis that, instead of cloning the entire output distributions on differing elements, clones only the portions of those distributions where they actually differ.
Specifically, it leverages a lemma from (Ye and Barg, 2017) to establish the existence of the following decomposition:
Theorem 6 (Corollary 3.4 in (Feldman et al., 2023a)).
Given any $\epsilon_{0}$-DP local randomizer $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and any $n+1$ inputs $x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}\in\mathbb{X}$, if $\mathbb{Y}$ is finite then there exists $p\in[0,1/(e^{\epsilon_{0}}+1)]$ and distributions $\mathcal{Q}_{1}^{0},\mathcal{Q}_{1}^{1},\mathcal{Q}_{1},\mathcal{Q}_{2},\dots,\mathcal{Q}_{n}$ such that
$\mathcal{R}(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\epsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\forall i\in[2,n],\,\mathcal{R}(x_{i})$ $=p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-2p)\mathcal{Q}_{i}.$
Such a decomposition is guaranteed to exist for any local randomizer. However, an error occurred in the construction of the reduction pair $P_{0}$ and $P_{1}$ based on this decomposition. Similar to the standard clone in Section 3.1, they define the following distribution $P_{0}(\epsilon_{0},p)$ and $P_{1}(\epsilon_{0},p)$: For any $p\in[0,1/(e^{\epsilon_{0}}+1)]$, let
$C\sim\text{Bin}(n-1,2p),\ A\sim\text{Bin}(C,1/2)$
and
$\Delta_{1}\sim\text{Bern}(e^{\epsilon_{0}}p),\quad\Delta_{2}\sim\text{Bin}(1-\Delta_{1},p/(1-e^{\epsilon_{0}}p)).$
Let
$P_{0}(\epsilon_{0},p)=(A+\Delta_{1},C-A+\Delta_{2})\quad\text{and}\quad P_{1}(\epsilon_{0},p)=(A+\Delta_{2},C-A+\Delta_{1}).$
They intended to prove that
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p),P_{1}(\epsilon_{0},p)\big{)},$
which serves as an upper bound for a specific local randomizer (different $\epsilon_{0}$-DP randomizers may have different values of $p$). Leveraging Lemma 7, they would then conclude the general upper bound for any $\epsilon_{0}$-DP local randomizer:
(1) $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},\frac{1}{e^{\epsilon_{0}}+1}),P_{1}(\epsilon_{0},\frac{1}{e^{\epsilon_{0}}+1})\big{)}.$
Lemma 7 (Lemma 5.1. in (Feldman et al., 2023a)).
For any $p,p^{\prime}\in[0,1]$ and $\varepsilon>0$, if $p<p^{\prime}$, then
$D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\,\|\,P_{1}(\varepsilon_{0},p)\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p^{\prime})\,\|\,P_{1}(\varepsilon_{0},p^{\prime})\big{)}$
Unfortunately, they encountered difficulty in constructing a post-processing function $f$ for this construction of $P_{0}$ and $P_{1}$. While they provided a function in the original paper, it was proven to be incorrect in the corrected revision (Feldman et al., 2023b). The issue arises from the fact that the “leftover” distribution of $x_{1}$ (i.e., $\mathcal{Q}_{1}$) is mixed with the “leftover” distribution of $x_{i}$ (i.e., $\mathcal{Q}_{i}$) in the above construction. In this case, the function $f$ does not know which distribution to sample from. This technical problem is fundamental and remains unsolved.
Although the corrected version was published on arXiv in October 2023, this error has been propagated in subsequent works (Wang et al., 2024; Chen et al., 2024; Yang et al., 2024b; Wang et al., 2025). For instance, the variation-ratio framework made significant efforts to design an algorithm to find the parameter $p$ for various specific local randomizers in the decomposition of Theorem 6. However, their work relies on the incorrect post-processing function $f$ presented in (Feldman et al., 2023a), which renders their results invalid.
Due to this fundamental difficulty, it is required that $x_{1}$ has no “leftover” distribution. In other words, each component of $\mathcal{R}(x_{1})$ must be distinguishable from the “leftover” distribution of $x_{i}$ for $i\geq 2$. In the above example, this necessitates a four-point-based construction for $P_{0}(\epsilon_{0},p,q)$ and $P_{1}(\epsilon_{0},p,q)$ (Feldman et al., 2023b):
$\mathcal{R}(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\mathcal{R}(x_{1}^{1})$ $=p\mathcal{Q}_{1}^{0}+e^{\epsilon_{0}}p\mathcal{Q}_{1}^{1}+(1-p-e^{\epsilon_{0}}p)\mathcal{Q}_{1},$
$\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=p\mathcal{Q}_{1}^{0}+p\mathcal{Q}_{1}^{1}+q\mathcal{Q}_{1}+(1-2p-q)\mathcal{Q}_{i}.$
This new decomposition proposed in the corrected version introduces additional challenges. First, for specific randomizers, computing a tight value of $q$ is nontrivial. Second, the monotonicity of $D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p,q),P_{1}(\epsilon_{0},p,q)\big{)}$ with respect to $p$ is not known. Consequently, we are unable to derive the desired conclusion—namely, a general upper bound applicable to any $\epsilon_{0}$-DP local randomizer, as stated in Formula (1). For the same reason, it remains unclear whether this decomposition necessarily yields tighter bounds than the standard clone decomposition. More critically, the new bound $D_{e^{\epsilon}}\big{(}P_{0}(\epsilon_{0},p,q),P_{1}(\epsilon_{0},p,q)\big{)}$ for a specific local randomizer lacks an efficient algorithm to compute.
4. General Clone and the Optimal Bounds
In this section, we formalize the general clone paradigm, which unifies and generalizes all decomposition methods for analyzing privacy amplification in the shuffle model. We then identify the optimal bounds achievable within this paradigm. The main results are summarized as follows:
  • Upper bound limitation: The general clone paradigm does not provide tighter bounds than the privacy blanket. In other words, its analytical capability is not inherently stronger than that of the privacy blanket framework.
  • Equivalence for specific randomizers: For any specific local randomizer, the optimal decomposition under the general clone paradigm is equivalent to the decomposition used in the privacy blanket framework.
The hierarchy of the bounds provided by the decomposition-based methods is shown in Figure 1.
Figure 1. Hierarchy among decompositions-based methods
4.1. Definition of general clone
Definition 0 (Decomposition in the General Clone Paradigm).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer. The general clone paradigm considers the following decomposition:
$\mathcal{R}(x_{1}^{0})$ $=\sum_{j=1}^{k}a_{j}Q_{1}^{j},$
$\mathcal{R}(x_{1}^{1})$ $=\sum_{j=1}^{k}b_{j}Q_{1}^{j},$
(2) $\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{k}c_{j}Q_{1}^{j}+\beta Q_{i},$
where $Q_{1}^{j}$ for $j=1,\dots,k$ and $Q_{i}$ for $i=2,\dots,n$ are probability distributions over $\mathbb{Y}$, and $a_{j},b_{j},c_{j},\beta$ are non-negative coefficients satisfying:
$\sum_{j=1}^{k}a_{j}=1,\quad\sum_{j=1}^{k}b_{j}=1,\quad\sum_{j=1}^{k}c_{j}+\beta=1.$
The general clone paradigm characterizes the general form of decompositions used in privacy amplification analysis. It directly subsumes the decomposition used in the standard clone framework. Although the decomposition defined by the privacy blanket appears structurally different, we will show that it naturally corresponds to a valid decomposition under the general clone paradigm.
When deriving the reduction pair from a general clone decomposition, an important constraint must be considered. Motivated by the failure of the stronger clone, the components $Q_{1}^{j}$ (shared across users) should not be mixed with the left-over distributions $Q_{i}$ (which are user-specific). This separation is essential to ensure the correctness and validity of the reduction.
Definition 0 (Reduction Pair of the General Clone).
Let $A_{0},A_{1}$, and $A_{2}$ be random variables on $\{1,2,\dots,k+1\}$, defined by the probabilities:
$\Pr[A_{0}=j]=\begin{cases}a_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases},\Pr[A_{1}=j]=\begin{cases}b_{j}&\text{for }j\leq k,\\ 0&\text{for }j=k+1.\end{cases}$
$\Pr[A_{2}=j]=\begin{cases}c_{j}&\text{for }j\leq k,\\ \beta&\text{for }j=k+1.\end{cases}$
The reduction pair $(P_{0}^{GC},P_{1}^{GC})$ (with “GC” denoting “General Clone”) is defined as the distributions of the histograms over $\{1,\dots,k+1\}$ generated by sampling:
  • One sample from $A_{0}$ for $P_{0}^{GC}$ (or $A_{1}$ for $P_{1}^{GC}$);
  • $n-1$ i.i.d. samples from $A_{2}$.
The output is the histogram vector $(n_{1},n_{2},\dots,n_{k},n_{k+1})$ indicating the counts of each index.
4.2. General clone is not stronger than blanket
Given a specific randomizer, we compare any decomposition within the general clone paradigm against the decomposition provided by the privacy blanket framework.
Theorem 3.
For every local randomizer, there is a post-processing function from $(P_{0}^{GC},P_{1}^{GC})$ to the $(P_{0}^{B},P_{1}^{B})$, where $(P_{0}^{GC},P_{1}^{GC})$ is the reduction pair given by the general clone paradigm equipped with any decomposition, and $(P_{0}^{B},P_{1}^{B})$ is the reduction pair given by the privacy blanket framework. Therefore,
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X_{0}),\mathcal{M}_{S}(X_{1}))\leq D_{e^{\epsilon}}(P_{0}^{B},P_{1}^{B})\leq D_{e^{\epsilon}}(P_{0}^{GC},P_{1}^{GC}).$
Proof.
The core idea of the proof is that the privacy blanket characterizes the maximal common part shared by all $\mathcal{R}(x_{i})$. The common part in any decomposition under the general clone paradigm cannot exceed that of the privacy blanket.
Recall the definition of $\omega(y)=\inf_{x\in\mathbb{X}}R(x)(y)/\gamma$, where $\gamma$ is a normalization factor. An important observation is that
$\forall y\in\mathbb{Y}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq\gamma\omega(y).$
It is because the formula (2) should always hold for all $x_{2},x_{3},\dots,x_{n}\in\mathbb{X}$, i.e., $\forall x\in\mathbb{X}:\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)\leq R(x)(y)$. Hence, it follows that $\sum_{j=1}^{k}c_{j}\leq\gamma$ and $1-\gamma\leq\beta$. It means that each $Q_{i}$ can be decomposed by
$Q_{i}=\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta}Q^{com}+(1-\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta})Q_{i}^{\prime},$
where $Q^{com},Q_{i}^{\prime}$ are two probability distributions, $Q^{com}$ is the common part of all $Q_{i}$.
The function $f_{2}$ shown in Algorithm 3 is the post-processing function satisfying $f_{2}(P_{0}^{GC})=P_{0}^{B},f_{2}(P_{1}^{GC})=P_{1}^{B}$. It behaves as follows: when encountering an index $i\in[1,k]$, it samples from the corresponding distribution $Q_{1}^{i}$; when encountering $i=k+1$, it samples from the common distribution $Q^{\text{com}}$ with a certain probability. Taken together, this behavior is equivalent to each user (except the first) sampling from the blanket distribution with probability $\gamma$. ∎
As a result, although both $D_{e^{\epsilon}}(P_{0}^{GC}||P_{1}^{GC})$ and $D_{e^{\epsilon}}(P_{0}^{B}||P_{1}^{B})$ serve as upper bounds on the privacy amplification in the shuffle model, the bound provided by the privacy blanket is always at least as tight as that of the general clone paradigm under any decomposition.
4.3. Blanket is “in” the General Clone
For every local randomizer, the general clone paradigm always admits a decomposition that is equivalent to the decomposition in the privacy blanket framework, where the components are single-point distributions $\mathbbm{1}_{y_{j}}$, with $y_{j}\in\mathbb{Y}$.
Theorem 4.
For every local randomizer, the following decomposition in the general clone paradigm is equivalent to the privacy blanket framework:
$\mathcal{R}(x_{1}^{0})$ $=\sum_{j=1}^{|\mathbb{Y}|}a_{j}\mathbbm{1}_{y_{j}}$
$\mathcal{R}(x_{1}^{1})$ $=\sum_{j=1}^{|\mathbb{Y}|}b_{j}\mathbbm{1}_{y_{j}}$
(3) $\forall i\in[2,n],\quad\mathcal{R}(x_{i})$ $=\sum_{j=1}^{|\mathbb{Y}|}c_{j}\mathbbm{1}_{y_{j}}+\beta Q_{i}.$
where $a_{j}=\mathcal{R}(x_{1}^{0})(y_{j})$, $b_{j}=\mathcal{R}(x_{1}^{1})(y_{j})$, and $c_{j}=\inf_{x\in\mathbb{X}}\mathcal{R}(x)(y_{j})$.
Proof.
It is straightforward to observe that $(P_{0}^{B},P_{1}^{B})$ and $(P_{0}^{GC},P_{1}^{GC})$ are essentially equivalent, differing only in some technical notations. ∎
We refer to the optimal decomposition of a local randomizer as the decomposition equivalent to that in the privacy blanket framework. For simplicity, we can merge some components to obtain an equally optimal decomposition. As an example, consider the $k$-Random Response mechanism.
Example 4.0.
Denote $\{1,2,\dots,k\}$ by $[k]$ and the uniform distribution on $[k]$ by $\mathcal{U}_{[k]}$. For any $k\in\mathbb{N}$ and $\varepsilon_{0}>0$, the $k$-randomized response mechanism $k\text{RR}:[k]\to[k]$ is defined as:
$k\text{RR}(x)=\begin{cases}x,&\text{with probability }\frac{e^{\varepsilon_{0}}-1}{e^{\varepsilon_{0}}+k-1},\\ y\sim\mathcal{U}_{[k]},&\text{with probability }\frac{k}{e^{\varepsilon_{0}}+k-1}.\end{cases}$
Its optimal decomposition is as follows:
$R(x_{1}^{0})$ $=e^{\epsilon_{0}}p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU,$
$R(x_{1}^{1})$ $=p\mathbbm{1}_{x_{1}^{0}}+e^{\epsilon_{0}}p\mathbbm{1}_{x_{1}^{1}}+qU,$
$\forall i\in[2,n]:R(x_{i})$ $=p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU+(e^{\epsilon_{0}}-1)p\mathbbm{1}_{x_{i}}.$
where $p=\frac{1}{e^{\epsilon_{0}}+k-1}$, $q=(k-2)p$, and $U$ is the uniform distribution over $[k]-\{x_{1}^{0},x_{1}^{1}\}$. This decomposition is also considered in the corrected version of the stronger clone (Feldman et al., 2023b).
Algorithm 3 Post-processing function from general clone paradigm to privacy blanket framework, $f_{2}$
$(n_{1},\dots,n_{k},n_{k+1})\in\{0,1,\dots,n\}^{k}$
$Y\leftarrow\emptyset$
for $i=1,\dots,k+1$ do
     for $count=1,2,\dots,n_{i}$ do
         if $i=k+1$ then
              $r\leftarrow\text{Bern}(\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta})$
              if $r=1$ then
                  $s\leftarrow_{Q^{com}}\mathbb{Y}$triangleright Sample from $Q^{com}$
                  $Y\leftarrow Y\cup\{s\}$
              end if
         else
              $s\leftarrow_{Q^{i}_{1}}\mathbb{Y}$triangleright Sample from $Q^{i}_{1}$
              $Y\leftarrow Y\cup\{s\}$
         end if
     end for
end for
return $(|Y|-1,Y)$
5. New Algorithm on Privacy Amplification
In this section, we revisit the previously overlooked concept of the privacy amplification random variable (PARV) within the privacy blanket framework. The PARV yields the exact amplification bound under the privacy blanket framework. However, the original paper did not provide a method for further simplification, resulting in only a loose closed-form bound and no approach for precise numerical computation. We address these limitations by proposing the generalized privacy amplification random variable, which resolves both issues. Our method achieves the best-known results under the general clone paradigm.
5.1. (Generalized) Privacy amplification random variable
Definition 0 (Privacy Amplification Random Variable (Balle et al., 2019)).
Suppose $W\sim\omega$ is a $\mathbb{Y}$-valued random variable sampled from the blanket. For any $\epsilon>0$ and $x,x^{\prime}\in\mathbb{X}$, the privacy amplification random variable is defined as
$L^{x,x^{\prime}}_{\epsilon}=\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\omega(W)}.$
Using PARV, Balle et al. derived the precise expression for $D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})$:
Lemma 2 (Lemma 5.3 in (Balle et al., 2019)).
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. Fix $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$. Suppose $L_{1},L_{2},\dots$ are i.i.d. copies of $L_{\varepsilon}^{x_{1}^{0},x_{1}^{1}}$ and $\gamma$ is defined as in Section 3.2. Then, we have:
$D_{e^{\epsilon}}$ $(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))$
(4) $\leq D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{\gamma n}\sum_{m=1}^{n}\binom{n}{m}\gamma^{m}(1-\gamma)^{n-m}\mathbb{E}\left[\sum_{i=1}^{m}L_{i}\right]_{+}.$
The bound above can also be expressed probabilistically as follows (Balle et al., 2019). Let $M\sim\text{Bin}(n,\gamma)$ be the random variable counting the number of users who sample from the blanket of $\mathcal{R}$. Formula (4) can be re-written as:
$D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{\gamma n}\mathbb{E}_{M\sim\text{Bin}(n,\gamma)}\left[\sum_{i=1}^{M}L_{i}\right]_{+},$
where we use the convention $\sum_{i=1}^{m}L_{i}=0$ when $m=0$.
Unfortunately, Balle et al. stopped at this point and did not pursue further simplification. In the following, we demonstrate how to improve upon the PARV to enable precise computation.
Definition 0 (Generalized Privacy Amplification Random Variable (GPARV)).
Define the generalized privacy amplification random variable as:
$\text{G}_{\epsilon}^{x,x^{\prime}}=\begin{cases}\frac{1}{\gamma}L_{\epsilon}^{x,x^{\prime}},&\text{w.p.}\ \gamma,\\ 0,&\text{w.p.}\ 1-\gamma.\end{cases}$
where $L_{\epsilon}^{x,x^{\prime}}$ is defined in Definition 1.
We can now restate Lemma 2 in a simplified form:
Theorem 4.
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. Fix $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$. Suppose $G_{1},G_{2},\dots$ are i.i.d. copies of $G_{\varepsilon}^{x_{1}^{0},x_{1}^{1}}$ and $\gamma$ is defined as in Section 3.2. Then, we have:
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))\leq D_{e^{\epsilon}}(P_{0}^{B}\|P_{1}^{B})=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
The GPARV has the following properties:
Property 2.
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be an $\varepsilon_{0}$-LDP local randomizer. For any $\varepsilon\geq 0$ and $x,x^{\prime}\in\mathbb{X}$, the generalized privacy amplification random variable $G=G_{\varepsilon}^{x,x^{\prime}}$ satisfies:
  1. (1)
    $\mathbb{E}[G]=1-e^{\epsilon}$,
  2. (2)
    $1-e^{\epsilon+\epsilon_{0}}\leq G\leq e^{\epsilon_{0}}-e^{\epsilon}$.
Proof.
The first property follows from direct computation:
$\mathbb{E}[G]=\gamma\mathbb{E}\left[\frac{1}{\gamma}L\right]=\mathbb{E}[L]=\mathbb{E}_{W\sim\omega}\left[\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\omega(W)}\right]=1-e^{\epsilon}.$
The second property is due to the $\epsilon_{0}$-DP property of $R$: $\forall x\in\mathbb{X},y\in\mathbb{Y}:1\leq\frac{R(x)(y)}{\gamma\omega(y)}\leq e^{\epsilon_{0}}$, so
$1-e^{\epsilon+\epsilon_{0}}\leq\frac{R(x)(W)-e^{\epsilon}R(x^{\prime})(W)}{\gamma\omega(W)}\leq e^{\epsilon_{0}}-e^{\epsilon}.\qed$
Remark 1.
In the original paper of the privacy blanket framework, a similar property of PARV was provided, but in a loose form (Balle et al., 2019). Specifically, they established that $\gamma(e^{-\epsilon_{0}}-e^{\epsilon+\epsilon_{0}})\leq L\leq\gamma(e^{\epsilon_{0}}-e^{\epsilon-\epsilon_{0}})$. However, the $\epsilon_{0}$-DP property of $R$ actually guarantees a tighter bound: $\gamma(1-e^{\epsilon+\epsilon_{0}})\leq L\leq\gamma(e^{\epsilon_{0}}-e^{\epsilon})$.
5.2. New Algorithm for Computing Privacy Amplification Upper Bounds
We present a new algorithm for computing the optimal privacy amplification bound under the general clone paradigm for any specific local randomizer, as described in Algorithm 4.
Overview.
First, the distribution of the generalized privacy amplification random variable (GPARV) $G_{\epsilon}^{x,x^{\prime}}$ is computed for a given local randomizer. Since most local randomizers exhibit input symmetry, the distribution of $G$ typically does not depend on the specific values of $x$ and $x^{\prime}$. We thus denote it simply by $G_{\epsilon}$. The distributions of $G_{\epsilon}$ for commonly used local randomizers are summarized in Table 1.
Given the distribution of $G_{\epsilon}$, our algorithm discretizes it to obtain $\bar{G}$, by rounding each value in $G$ up to the nearest larger multiple of a discretization interval length $l$:
$\text{Round}(x)=l\cdot\lceil\frac{x}{l}\rceil.$
Next, the algorithm computes the $n$-fold convolution of $\bar{G}$, denoted $\bar{G}^{*n}$, using the classical Fast Fourier Transform method:
$\bar{G}^{*n}=\text{FFT}^{-1}\left((\text{FFT}(\bar{G}))^{\odot n}\right),$
where $\odot n$ represents the element-wise exponentiation by $n$.
Finally, the algorithm evaluates the integral
[…]
and outputs $\frac{I}{n}$ as the upper bound on privacy amplification.
Correctness.
The output of the algorithm is guaranteed to upper bound the true value, since the discretized distribution $\bar{G}$ stochastically dominates $G$:
$\frac{1}{n}\mathbb{E}[\bar{G}^{*n}]_{+}\geq\frac{1}{n}\mathbb{E}[G^{*n}]_{+}.$
Time Complexity and Error Analysis.
Because $G$ is supported on the interval $[1-e^{\epsilon+\epsilon_{0}},e^{\epsilon_{0}}-e^{\epsilon}]$ (Property 2), the discretization step requires $O\left(\frac{e^{\epsilon_{0}}}{l}\right)$ operations. For most frequency oracles used on categorical data, $G$ takes values on at most five points (see Table 1), resulting in $O(1)$ discretization time.
The discretization introduces a bounded error. An intuitive analysis is as follows:
$\mathbb{E}[\bar{G}^{*n}]_{+}-nl\leq\mathbb{E}[G^{*n}]_{+}\leq\mathbb{E}[\bar{G}^{*n}]_{+}.$
The FFT computation runs in $O\left(\frac{n}{l}\log\left(\frac{n}{l}\right)\right)=\widetilde{O}(\frac{n}{l})$ time. Choosing $l=O\left(\frac{1}{n}\right)$ ensures an $O(1)$ additive error in total $\widetilde{O}(n^{2})$ time.
More precise analysis can be done:
$\mathbb{E}[\bar{G}^{*n}]_{+}-nl\cdot\Pr\left[[G^{*n}]_{+}>-nl\right]\leq\mathbb{E}[G^{*n}]_{+}\leq\mathbb{E}[\bar{G}^{*n}]_{+}.$
By Hoeffding’s inequality,
$\Pr\left[[G^{*n}]_{+}>-nl\right]\leq\exp\left(-\left(\frac{2a^{2}}{b^{2}}-l\right)n\right),$
where $a=-\mathbb{E}[G]=e^{\epsilon}-1$ and $b=(e^{\epsilon_{0}}-1)(e^{\epsilon}+1)$. This shows that the error decays exponentially in $n$ as long as $l\leq\frac{2a^{2}}{b^{2}}$. Empirical evaluations confirm that $l=O(1)$ is sufficiently accurate in practice (see Section 7), and the overall FFT runtime becomes $O(n\log n)$.
Comparison with Existing Numerical Methods.
The numerical computation of privacy amplification in the shuffle model has been studied since the introduction of the clone paradigm by Feldman et al. (Feldman et al., 2022). Prior numerical algorithms can only handle three-point decompositions, such as computing $D_{e^{\epsilon}}(P_{0}^{C}\|P_{1}^{C})$ (Feldman et al., 2022; Wang et al., 2024; Koskela et al., 2021). However, these techniques do not generalize to decompositions involving more than three points, which are essential for obtaining tight bounds from optimal decompositions.
In contrast, our FFT-based algorithm supports the optimal decomposition of any local randomizer, enabling tighter and more accurate privacy amplification bounds. Furthermore, our method is not only more general but also simpler and significantly faster than existing numerical algorithms (see Section 7).
Table 1. The probability distribution of GPARV for common LDP protocols
$1-e^{\epsilon+\epsilon_{0}}$ $e^{\epsilon_{0}}-e^{\epsilon+\epsilon_{0}}$ $1-e^{\epsilon}$ 0 $e^{\epsilon_{0}}-e^{\epsilon}$
$k$-RR (Warner, 1965) $\frac{1}{e^{\epsilon_{0}}+k-1}$ 0 $\frac{k-2}{e^{\epsilon_{0}}+k-1}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+k-1}$ $\frac{1}{e^{\epsilon_{0}}+k-1}$
BLH (Wang et al., 2017) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+1}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
RAPPOR (Erlingsson et al., 2014) $\frac{1}{(e^{\epsilon_{0}/2}+1)^{2}}$ $\frac{1}{e^{\epsilon_{0}/2}(e^{\epsilon_{0}/2}+1)^{2}}$ $\frac{e^{\epsilon_{0}/2}}{(e^{\epsilon_{0}/2}+1)^{2}}$ $1-e^{-\epsilon_{0}/2}$ $\frac{1}{(e^{\epsilon_{0}/2}+1)^{2}}$
OUE (Wang et al., 2017) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2e^{\epsilon_{0}}(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2}(1-e^{-\epsilon_{0}})$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
HR (Acharya et al., 2019) $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$ $\frac{e^{\epsilon_{0}}-1}{e^{\epsilon_{0}}+1}$ $\frac{1}{2(e^{\epsilon_{0}}+1)}$
Algorithm 4 Calculate the optimal privacy amplification bound via FFT
the distritbuion density $G$ of $G_{\epsilon}$, number of users $n$, discretisation interval length $l$
$\bar{G}=$ Discretize($G$,$l$) $\triangleright$ Round every point of $G$ to $nl,n\in\mathbb{Z}$
$\bar{G}^{*}=\text{FFT}^{-1}((\text{FFT}(\bar{G}))^{\odot n})$$\triangleright$ Compute the $n$-fold convolution
[…]
return $\frac{I}{n}$
5.3. New Amplification Lower Bounds
Our generalized privacy amplification random variable (GPARV) also facilitates the computation of lower bounds for privacy amplification via shuffling. These bounds help to demonstrate the tightness of the upper bounds.
An upper bound refers to the existence of a value $\delta_{u}$ such that, for a given local randomizer with specified $\epsilon_{0},\epsilon,n$, and for any two neighboring input datasets $X^{0}$ and $X^{1}$, we have:
$D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}\leq\delta_{u}.$
For a given local randomizer with specified $\epsilon_{0},\epsilon,n$, we can construct two neighboring datasets $X^{0}$ and $X^{1}$, and compute:
$\delta_{l}=D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)},$
which serves as the lower bound for this amplification.
In previous studies, the common approach to selecting $X^{0}$ and $X^{1}$ is to set $x_{2}=x_{3}=\dots=x_{n}$ such that $x_{1}^{0},x_{1}^{1},x_{2}$ are mutually distinct (Feldman et al., 2023a; Wang et al., 2024). The following theorem provides an efficient method for computing $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$ under this setting.
Theorem 5.
Let $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ be a local randomizer, and let $\mathcal{M}_{S}=\mathcal{S}\circ\mathcal{R}^{n}$ be the shuffling of $\mathcal{R}$. Fix $\varepsilon\geq 0$ and inputs $X^{0}\simeq X^{1}$ with $x_{1}^{0}\neq x_{1}^{1}$ and $x_{2}=x_{3}=\dots=x_{n}$. Define a random variable $G=\frac{\mathcal{R}(x_{1}^{0})(y)-e^{\epsilon}\mathcal{R}(x_{1}^{1})(y)}{\mathcal{R}(x_{2})(y)}$, where $y\sim\mathcal{R}(x_{2})$. Suppose $G_{1},G_{2},\dots$ are i.i.d. copies of $G$. Then, we have the following:
$D_{e^{\epsilon}}(\mathcal{M}_{S}(X^{0})\|\mathcal{M}_{S}(X^{1}))=\frac{1}{n}\mathbb{E}\left[\sum_{i=1}^{n}G_{i}\right]_{+}.$
We provide the proof in Appendix B. The result indicates that, for $X^{0}$ and $X^{1}$ satisfying $x_{2}=x_{3}=\dots=x_{n}$, we can efficiently compute $D_{e^{\epsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}$ using FFT.
6. Joint Composition of LDP
In this section, we analyze the joint composition of multiple LDP local randomizers $\mathcal{R}_{1},\mathcal{R}_{2},\dots,\mathcal{R}_{m}$. Let $\mathcal{R}_{i}:\mathbb{X}^{i}\to\mathbb{Y}^{i}$ denote the local randomizer for the $i$-th component. The joint composition of these randomizers is defined as follows:
$\mathcal{R}^{\times}_{[1:m]}:\mathbb{X}^{1}\times\mathbb{X}^{2}\times\dots\times\mathbb{X}^{m}\to\mathbb{Y}^{1}\times\mathbb{Y}^{2}\times\dots\times\mathbb{Y}^{m},$
$\mathcal{R}^{\times}_{[1:m]}(x_{1},x_{2},\dots,x_{m})=(\mathcal{R}_{1}(x_{1}),\mathcal{R}_{2}(x_{2}),\dots,\mathcal{R}_{m}(x_{m})).$
We assume that any two input vectors $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{m})$ and $\boldsymbol{x}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\dots,x_{m}^{\prime})$ are neighboring. In practice, it is common to use the joint composition of $m$ mechanisms, each satisfying $\left(\frac{\epsilon}{m}\right)$-LDP, to achieve overall $\epsilon$-LDP.
Theorem 1 (Composition Theorem).
If $\mathcal{R}_{i}$ satisfies $\epsilon_{i}$-LDP, then $\mathcal{R}^{\times}_{[1:m]}$ satisfies $\sum_{i=1}^{m}\epsilon_{i}$-LDP.
When applied to LDP protocols resulting from joint composition, the clone paradigm provides particularly loose decompositions. This is due to the presence of many intermediate states in the joint probability distribution after the composition. For instance, most local randomizers have probability ratios between any two inputs that belong to the set $\{e^{-\epsilon},1,e^{\epsilon}\}$. However, after a $m$-fold joint composition, the ratio of the joint probability distribution may belong to $\{e^{\frac{i}{m}\epsilon}|i\in[-m,m]\cap\mathbb{Z}\}$. Under these conditions, the common part in the clone paradigm deviates significantly from the actual privacy blanket.
To compute the privacy amplification induced by the privacy blanket, we only need to compute the GPARV of $\mathcal{R}^{\times}_{[1:m]}$, which is equivalent to finding the optimal decomposition of $\mathcal{R}^{\times}_{[1:m]}$. Fortunately, the optimal decomposition of a joint composition mechanism is simply the Cartesian product of the optimal decompositions of each individual LDP component.
As an example, consider the joint composition of two $k$-RR mechanisms acting on $[k]\times[k]$. The optimal decomposition is given below:
Example 6.0 (Optimal Decomposition of Two-Joint $k$-RR).
$\mathcal{R}(a_{1}^{0},b_{1}^{0})=\left(e^{\epsilon_{0}/2}p\mathbbm{1}_{a_{1}^{0}}+p\mathbbm{1}_{a_{1}^{1}}+qU\right)\times\left(e^{\epsilon_{0}/2}p\mathbbm{1}_{b_{1}^{0}}+p\mathbbm{1}_{b_{1}^{1}}+qU\right),$
$\mathcal{R}(a_{1}^{1},b_{1}^{1})=\left(p\mathbbm{1}_{a_{1}^{0}}+e^{\epsilon_{0}/2}p\mathbbm{1}_{a_{1}^{1}}+qU\right)\times\left(p\mathbbm{1}_{b_{1}^{0}}+e^{\epsilon_{0}/2}p\mathbbm{1}_{b_{1}^{1}}+qU\right),$
$\forall i\in[2,n],\ \mathcal{R}(a_{i},b_{i})=\left(p\mathbbm{1}_{a_{1}^{0}}+p\mathbbm{1}_{a_{1}^{1}}+qU+(1-2p-q)\mathbbm{1}_{a_{i}}\right)$
×$\quad\times\left(p\mathbbm{1}_{b_{1}^{0}}+p\mathbbm{1}_{b_{1}^{1}}+qU+(1-2p-q)\mathbbm{1}_{b_{i}}\right)$
where $p=\frac{1}{e^{\epsilon_{0}/2}+k-1}$, $q=\frac{k-2}{e^{\epsilon_{0}/2}+k-1}$, and $U$ is the uniform distribution over $[k]-\{a_{1}^{0},a_{1}^{1}\}$. The Cartesian product of two probability distributions $P,Q$ is defined as $(P\times Q)(a,b)=P(a)\cdot Q(b)$.
7. Numerical Experiments
In this section, we conduct experimental evaluations of our FFT-based numerical algorithm proposed in this paper.
We compare the optimal bounds for specific local randomizers under the general clone paradigm with existing bounds. The baselines include two bounds from the privacy blanket framework, derived using Hoeffding’s and Bennett’s inequalities, respectively (Balle et al., 2019), as well as the numerical bounds from the standard clone paradigm (Feldman et al., 2022). We utilized the open-source implementations released by the respective papers.
We present results for four commonly used local randomizers: $k$-Randomized Response (Warner, 1965) ($k=10$), Binary Local Hash (Wang et al., 2017), RAPPOR (Erlingsson et al., 2014), and Optimized Unary Encoding (OUE) (Wang et al., 2017). The experimental results are shown in Figure 2 and Figure 3. The lower bounds are computed using the method described in Section 5.3. In the legends of the plots, “$k$-joint” refers to the joint composition of $k$ local randomizers, each satisfying $\frac{\epsilon_{0}}{k}$-LDP. In all experiments, the discretization interval length $l$ of our algorithm is set as $\frac{e^{\epsilon_{0}}-1}{1200}$.
The experimental results show that our upper bounds consistently outperform existing methods. Furthermore, the gap between our upper and lower bounds is generally small, indicating the tightness and reliability of the computed results. This also validates that our choice of discretization interval length $l$ offers sufficient precision for practical use.
In addition to its accuracy, our algorithm is highly efficient: it generates a full curve in approximately 30 seconds, whereas the numerical algorithm used for the standard clone requires around 5 minutes. It is worth noting that the numerical generic bound of the standard clone is weaker than the specific one provided by the privacy blanket using Bennett’s inequality in our experiments.
Our results show that computing bounds specific to joint compositions leads to significantly tighter amplification bounds compared to generic methods. This highlights the importance and advantage of our algorithm in accurately analyzing privacy in practical multi-attribute settings. Furthermore, we observe that, under a fixed total privacy budget $\epsilon_{0}$, the privacy amplification effect achieved through shuffling becomes stronger as the number of composed randomizers $k$ increases.
10-RR, low $\epsilon_{0}$
10-RR, large $\epsilon_{0}$
BLH, low $\epsilon_{0}$
BLH, large $\epsilon_{0}$
Figure 2. Experimental Results: RR and BLH
RAPPOR, low $\epsilon_{0}$
RAPPOR, large $\epsilon_{0}$
OUE, low $\epsilon_{0}$
OUE, large $\epsilon_{0}$
Figure 3. Experimental Results: RAPPOR and OUE
8. Discussion: Beyond the General Clone
In this work, we develop an efficient algorithm to compute the best-known privacy amplification bounds within the general clone framework, which encompasses all possible decompositions. A natural and important question arises: Can we achieve tighter bounds than those provided by the general clone? Addressing this question requires stepping beyond decomposition methods.
A promising direction is to identify the most vulnerable neighboring dataset pair $(X^{0},X^{1})$ such that $D_{e^{\epsilon}}\left(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\right)$ is maximized among all neighboring pairs of size $n$. If, for a local randomizer $\mathcal{R}$, one can prove that a specific pair $(X^{0}_{v},X^{1}_{v})$ consistently maximizes the divergence for every $\epsilon$, then this would yield the exact privacy amplification bound for $\mathcal{R}$ under shuffling.
For many local randomizers, a plausible candidate for the most vulnerable dataset pair is
$X^{0}=(x_{1}^{0},x_{2},x_{2},\dots,x_{2}),\quad X^{1}=(x_{1}^{1},x_{2},x_{2},\dots,x_{2}),$
where $x_{1}^{0},x_{1}^{1},x_{2}\in\mathbb{X}$ are mutually distinct. This construction is also used in Section 5.3 for computing lower bounds of privacy amplification.
This conjecture is motivated by two observations. First, to maximize distinguishability, the set of inputs $\{x_{i}\mid i=2,3,\dots,n\}$ should exclude both $x_{1}^{0}$ and $x_{1}^{1}$, ensuring that the outputs are not easily confounded. Second, having unified inputs among the remaining users simplifies the inference of their output contributions, thereby potentially increasing the overall distinguishability between the shuffled outputs of $X^{0}$ and $X^{1}$.
Despite its intuitive appeal, this conjecture currently lacks a formal proof. Developing tools to rigorously establish the most vulnerable neighboring pair remains an open problem and a valuable direction for future research.
9. Conclusion
In this work, we propose the general clone framework, which encompasses all decomposition methods, and identify the optimal bounds within the general clone framework. We also present an efficient algorithm for numerically computing these optimal bounds. With these results, we achieve the best-known bounds. Experiments demonstrate the tightness of our analysis. Additionally, we explore the joint composition of LDP protocols in the shuffle model for the first time.
We hope that this work contributes to both the practical deployment and the theoretical advancement of the shuffle model in differential privacy.
Appendix A Misunderstanding of the Privacy Blanket Framework
In this section, we discuss and correct a misunderstanding in the literature regarding the privacy blanket framework (Koskela et al., 2021).
The original paper on the privacy blanket framework demonstrates its usage with the example of $k$-Random Response (see Definition 5) (Balle et al., 2019). The privacy blanket distribution for $k$-RR is simply the uniform distribution over $[k]$. Let $N_{0}$ and $N_{1}$ denote the counts of $x_{1}^{0}$ and $x_{1}^{1}$ appearing in $Y_{|\boldsymbol{b}|}$ (see Definition 4), respectively. When $b_{1}=1$, it is easy to see that
$\Pr[P_{0}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]=\Pr[P_{1}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0].$
When $b_{1}=0$, the first user commits their true value, i.e., $y_{1}^{b}=x_{1}^{b}$. In this case, using combinatorial analysis, we can derive the following (Balle et al., 2019):
$\forall Y_{|\boldsymbol{b}|}:\quad\frac{\Pr[P_{0}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]}{\Pr[P_{1}^{B}(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\mid b_{1}=0]}=\frac{N_{0}}{N_{1}}.$
The mistake in (Koskela et al., 2021) lies in their characterization of the joint distribution of $N_{0}$ and $N_{1}$:
$N_{0}\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon}+k-1})+1\text{ and }N_{1}\sim\text{Bin}(n-1,\frac{1}{e^{\epsilon}+k-1})$
under the condition $b_{1}=0$ and $(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})\sim P_{0}^{B}$. In reality, $N_{0}$ and $N_{1}$ are not independent, as also noted in (Luo et al., 2022). The correct characterization is as follows: let $N\sim\text{Bin}(n-1,\frac{k}{e^{\epsilon_{0}}+k-1})$, then $N_{0}\sim\text{Bin}(N,\frac{1}{k})+1$, and $N_{1}\sim\text{Bin}(N-N_{0},\frac{1}{k-1})$.
This mistake affects both the theoretical analysis and the experimental results reported in (Koskela et al., 2021).
Appendix B Proof of Theorem 5
Proof.
The proof of this theorem follows similar lines as the proof of Lemma 12 in (Balle et al., 2019).
Define random variables $Y_{1}^{b}\sim\mathcal{R}(x_{1}^{b})$ and $W_{i}\sim\mathcal{R}(x_{2}),i=1,2,\dots,n-1$. $\mathbf{W}_{n-1}=\{W_{1},W_{2},\dots,W_{n-1}\}$. Let $\vec{y}\in\mathbb{Y}^{n}$ be a tuple of elements from $\mathbb{Y}$ and $Y\in\mathbb{N}_{n}^{\mathbb{Y}}$ be the corresponding multiset of entries. Then we have
[$\mathbb{P}[\{Y_{1}^{b}\}\cup\mathbf{W}_{n-1}=Y]=\frac{1}{n!}\sum_{\sigma}\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right],$
where $\sigma$ ranges over all permutations of $[n]$ and $\vec{y}_{\sigma}=(y_{\sigma(1)},\ldots,y_{\sigma(n)})$. We also have
$\mathbb{P}\left[(Y_{1}^{b},W_{1},\ldots,W_{n-1})=\vec{y}_{\sigma}\right]=\mathcal{R}(x_{1}^{b})(y_{\sigma(1)})\prod_{i=2}^{n}\mathcal{R}(x_{2})(y_{\sigma(i)}).$
Summing this expression over all permutations $\sigma$ and factoring out the product of the $\mathcal{R}(x_{2})$’s yields:
$\frac{1}{n!}\sum_{\sigma}\mathcal{R}(x_{1}^{b})(y_{\sigma(1)})\mathcal{R}(x_{2})(y_{\sigma(2)})\cdots\mathcal{R}(x_{2})(y_{\sigma(n)})$
$=\left(\prod_{i=1}^{n}\mathcal{R}(x_{2})(y_{i})\right)\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{b})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right)$
$=\mathbb{P}[\mathbf{W}_{n}=Y]\cdot\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{b})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}.$
Now we can plug these observation into the definition of $\mathcal{D}_{e^{\varepsilon}}$ and complete the proof as
$\mathcal{D}_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big{)}=\mathcal{D}_{e^{\varepsilon}}(\{Y_{1}^{0}\}\cup W_{n-1}\parallel\{Y_{1}^{1}\}\cup W_{n-1})$
$=\int_{\mathbb{N}_{n}^{\mathbb{Y}}}\left[\mathbb{P}[\{Y_{1}^{0}\}\cup W_{n-1}=Y]-e^{\varepsilon}\mathbb{P}[\{Y_{1}^{1}\}\cup W_{n-1}=Y]\right]_{+}$
$=\int_{\mathbb{N}_{n}^{\mathbb{Y}}}\mathbb{P}[W_{n}=Y]\left[\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{0})(y_{i})-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\frac{\mathcal{R}(x_{1}^{0})(y_{i})-e^{\varepsilon}\mathcal{R}(x_{1}^{1})(y_{i})}{\mathcal{R}(x_{2})(y_{i})}\right]_{+}$
$=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}G_{i}\right]_{+}.\qed$